Entries for Factory Costs and Jobs Completed Old School Publishing Inc. began printing operations on January 1. Jobs 301 and 302 were completed during the month, and all costs applicable to them were recorded on the related cost sheets. Jobs 303 and 304 are still in process at the end of the month, and all applicable costs except factory overhead have been recorded on the related cost sheets. In addition to the materials and labor charged directly to the jobs, $3,080 of indirect materials and $37,280 of indirect labor were used during the month. The cost sheets for the four jobs entering production during the month are as follows, in summary form: Job 301 Job 302 Direct materials $31,880 Direct materials $14,900 Direct labor 12,200 Direct labor 6,300 Factory overhead 6,710 Factory overhead 3,465 Total $50,790 Total $24,665 Job 303 Job 304 Direct materials $46,400 Direct materials $9,180 Direct labor 13,500 Direct labor 1,800 Factory overhead — Factory overhead — Journalize the summary entry to record each of the following operations for January (one entry for each operation): a. Direct and indirect materials used. For a compound transaction, if an amount box does not require an entry, leave it blank. b. Direct and indirect labor used. For a compound transaction, if an amount box does not require an entry, leave it blank. c. Factory overhead applied to all four jobs (a single overhead rate is used based on direct labor cost). d. Completion of Jobs 301 and 302.
Answer #1

Ans- Journal Entries

Date Account Title and Explanation Debit Credit
1)- Work in progress $102,360
Factory Overhead $3,080
Material $105,440
(To record direct and indirect material used)
2)- Work in progress $33,800
Factory Overhead $37,280
Wages Payable $71,080
(To record direct and indirect labor used)
3)- Work in progress $18,590
Factory Overhead ($33,800*55%) $18,590
(To record factory overhead applied)
4)- Finished Goods ($50,790+ $24,665) $75,455
Work in progress $75,455
(To record completion of Job 301 and Job 302)

Predetermined overhead rate= Job 301-$6,710/ $12,200*100= 55% of direct labor cost

Job-302-$3,465/$6,300*100= 55% of direct labor cost

Factory Overhead applied to Job 303- $13,500*55%= $7,425

Factory Overhead applied to Job 304- $1,800*55%= 990

Kindly give thumbsup if u like my answer ....Thanks!!!